can you please help with this question? I am so lost with it.
the odds for even A are 5 to 7.
a) what is the probability that event A happens?
b) what is the probability that event A doeant happen?
c) what are the odds against A happening?

thank you thank you so much for your help!

Answer by ewatrrr(24785)


Hi
the odds for event A are 5 to 7.
a) what is the probability that event A happens? P = 5/12
b) what is the probability that event A doesnt happen? P = 1-5/12 = 7/12
c) what are the odds against A happening? 7 to 5